# Rubrics as Rewards: Reinforcement Learning Beyond Verifiable Domains

### v1

@@ Abstract
Extending Reinforcement Learning with Verifiable Rewards (RLVR) to real-world tasks often requires balancing objective and subjective evaluation criteria. However, many such tasks lack a single, unambiguous ground truth—making it difficult to define reliable reward signals for post-training language models. While traditional preference-based methods offer a workaround, they rely on opaque reward functions that are difficult to interpret and prone to spurious correlations. We introduce Rubrics as Rewards (RaR), a framework that uses structured, checklist-style rubrics as interpretable reward signals for on-policy training with GRPO. Our best RaR method yields up to a 28% relative improvement on HealthBench-1k compared to simple Likert-based approaches, while matching or surpassing the performance of reward signals derived from expert-written references. By treating rubrics as structured reward signals, we show that RaR enables smaller-scale judge models to better align with human preferences and sustain robust performance across model scales.
Rubrics as Rewards:
Reinforcement Learning Beyond Verifiable Domains
Anisha Gunjal   Anthony Wangthanks: Work done while at Scale AI.   Elaine Lau Vaskar NathBing LiuSean Hendryx Scale AI anisha.gunjal@scale.com
@@ 1 Introduction
Reinforcement Learning with Verifiable Rewards (RLVR) has enabled large language models to elicit complex reasoning on tasks with verifiable outcomes, most prominently in mathematics and coding, where a reward model can be replaced with a scoring function and final answers are easily verifiable or test cases provide automated mechanisms for evaluation Lambert et al. (2024); Guo et al. (2025a); Cui et al. (2025). However, many real-world tasks lack such explicit verifiable answers, leaving models without direct reward feedback. In practice, researchers often turn to RLHF via preference ranking—collecting human judgments over pairs or lists of model outputs to fill this gap. While preference-based reward models can bootstrap performance, they tend to overfit superficial artifacts (e.g. response length, formatting quirks, annotator biases) Singhal et al. (2023); Wang et al. (2024); Chen et al. (2024b); Ye et al. (2024); Gudibande et al. (2023) and require large volumes of pairwise comparisons Ouyang et al. (2022), making preference-based reward models brittle and costly.
To address this gap, we introduce Rubrics as Rewards (RaR) for on-policy Reinforcement Learning that treats structured criteria or rubrics as the core reward mechanism. We use rubrics as checklists Arora et al. (2025); Sirdeshmukh et al. (2025) composed of independent subgoals allowing for automatable feedback aligned with expert intent. By decomposing “what makes a good response” into tangible, human-interpretable criteria, rubrics offer a middle ground between binary correctness signals and coarse preference rankings.
@@
Previous works train generative reward models that learn to evaluate reasoning or final outputs with interpretable scores Chen et al. (2025); Whitehouse et al. (2025); Anugraha et al. (2025); Guo et al. (2025b), and some have even used a model’s internal confidence estimates as a proxy for reward Zhao et al. (2025). Concurrently, recent efforts have extended verifiable datasets beyond STEM domains, broadening the applicability of RLVR methods to a wider range of tasks Su et al. (2025b); Ma et al. (2025). However, developing a general-purpose approach for specifying reliable and scalable rewards remains an open challenge, particularly in settings where no single ground truth exists, or where both subjective and objective criteria must be considered to evaluate response quality. The Rubrics as Rewards strategy offers a flexible solution by repurposing structured evaluation criteria into multi-dimensional reward signals. Figure 1 illustrates our approach for generating rubrics and leveraging them as reward signals for on-policy training.
Our key contributions are as follows: (i) We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning framework that uses checklist-style rubrics to supervise multi-criteria tasks, enabling stable training and improved performance in both reasoning and real-world domains. (ii) We apply our rubric generation approach to the medicine and science domains, producing two training datasets—RaR-Medicine-20k and RaR-Science-20k [^1] The datasets will be released soon.. (iii) We show that models trained with RaR match or outperform strong baselines, achieving notable gains in accuracy across diverse domains. (iv) By using rubrics as structured reward signals, RaR enables smaller judge models to achieve better alignment with human preferences and maintains robust performance across different model scales.
@@ 2.1 Problem Formulation
Let $x$ denote an input prompt and $\hat{y}\sim\pi_{\theta}(\cdot\mid x)$ be a sampled response from a model parameterized by $\theta$. In domains without single ground-truth answers or automatic correctness signals, we define a structured reward function using prompt-specific rubric criteria.
Each prompt $x$ is associated with a set of $k$ rubric items $\{(w_{j},c_{j})\}_{j=1}^{k}$, where $w_{j}\in\mathbb{R}$ denotes the weight of criterion $j$, and $c_{j}:(x,\hat{y})\mapsto\{0,1\}$ is a binary correctness function that indicates whether the response $\hat{y}$ satisfies that criterion given the prompt. The final normalized scalar reward is computed as:
@@
This normalization ensures reward values are consistent across prompts with varying numbers and weights of rubric items.
Explicit Rubric Aggregation.
Each criterion is evaluated independently using a generative reward model, and the reward is computed using Equation 1. While we currently assume binary values for $c_{j}$, this framework supports extension to continuous-valued judgments.
Implicit Rubric Aggregation.
An alternative approach passes all rubric descriptions and weights to a generative reward model, which internally computes a scalar reward:
$r_{\text{implicit}}(x,\hat{y})=f_{\phi}(x,\hat{y},\{(w_{j},d_{j})\}_{j=1}^{k})$ (2)
Here, $f_{\phi}$ denotes an LLM-based judge that takes as input the prompt $x$, the generated response $\hat{y}$, and a set of weighted rubric criteria. Each pair $(w_{j},d_{j})$ consists of a scalar or categorical weights $w_{j}\in\mathbb{R}$ and a rubric criterion $d_{j}$ describing a specific aspect of response quality. This formulation delegates the aggregation of evaluation criteria to the model itself, allowing it to compute a holistic reward score.
In practice, we use an LLM-as-judge with separate prompt templates containing prompt, response, and rubric items for explicit and implicit aggregation. The prompts used in our experiments are detailed in Appendix E.
@@ 2.2 Generalization of RLVR with Rubrics as Rewards
Rubric-based reward functions thus generalize RLVR by enabling multi-dimensional supervision, flexible weighting across criteria, and the incorporation of both objective and subjective aspects of response quality.
This formalization highlights that RLVR can be seen as a restricted instance of rubric-guided RL with a single essential criterion. In contrast, rubric-based rewards further enable structured supervision in settings where correctness is multifaceted and may not be strictly verifiable.
@@ 3.1 Desiderata
We synthesize prompt-specific rubrics guided by a set of design principles. Each rubric defines the criteria for a high-quality response and enables human-interpretable supervision. We use large language models (LLMs) as expert proxies to generate these rubrics while ensuring adherence to the following desiderata:
Grounded in Expert-guidance
Reference answers produced by human experts or stronger LLMs serve as proxies for expert guidance. Rubrics are grounded in these references to capture the key facts, reasoning steps, and conclusions necessary for correctness.
Comprehensive coverage
Rubrics are designed to span multiple quality dimensions, including factual accuracy, logical structure, completeness, style, and common pitfalls. Negative “pitfall” criteria help penalize frequent or high-risk errors. The total number of rubric items per rubric (typically 7–20) is chosen to balance coverage with signal sparsity.
Semantic weighting
Each criterion is labeled with a categorical importance level (e.g., Essential, Important, Optional, Pitfall) that reflects its relative priority in the final reward. This supports interpretable reward aggregation. Future work may explore numeric or learned weighting for greater flexibility.
Self-contained evaluation
Rubrics are written to be independently actionable-each item can be evaluated in isolation by either human annotators or LLM-based judges, without requiring external context or domain knowledge.
3.2 Synthesis
We apply the above framework on real world science and Medicine reasoning domains. For each domain, the prompt (included in Appendix H) instructs the LLM to generate 7–20 rubric items based on the complexity of the input question. Each item is assigned a categorical weight (e.g., Essential Criteria, Important Criteria) to determine its importance to a correct answer. The rubrics are designed to be fully self-contained which means that non-expert readers should be able to evaluate response quality using only the rubric. The evaluation outcome can be directly used as reward functions during reinforcement learning.
We generate synthetic rubrics using OpenAI’s o3-mini and gpt-4o models OpenAI o3-mini (2025); Jaech et al. (2024); Hurst et al. (2024) as described in Section 3. The RaR-Medical-20k and RaR-Science-20k are released for public use. For both domains, we use the reference answers provided in the datasets as a proxy for human expertise while generating synthetic rubrics. These rubric sets are then used to supervise the training of smaller models through either explicit or implicit reward aggregation (Section 2.1).
@@ 4 Experiments
Domain $\rightarrow$ Medicine Science
Method/Baseline $\downarrow$ HealthBench-1k GPQA_Diamond
Qwen2.5-7b 0.0818 0.3030 $\pm$ 0.0286
Qwen2.5-7b-Instruct 0.2359 0.3598 $\pm$ 0.0077
Simple-Likert 0.2489 0.3409 $\pm$ 0.0104
Reference-Likert 0.3155 0.3775 $\pm$ 0.0350
Predefined RaR 0.2472 0.3485 $\pm$ 0.0365
RaR-Explicit (o3-mini rubrics) 0.2559 0.3333 $\pm$ 0.0504
RaR-Explicit (GPT-4o rubrics) 0.2979 0.3030 $\pm$ 0.0197
RaR-Implicit (o3-mini rubrics) 0.3107 0.3864 $\pm$ 0.0407
RaR-Implicit (GPT-4o rubrics) 0.3194 0.3662 $\pm$ 0.0191
Table 1: Performance of baseline and RaR (Rubrics as Rewards) methods on medicine and science domains. For HealthBench-1k, we report the overall score; for GPQA_Diamond, we report mean accuracy over 4 runs with standard deviation. All experiments use gpt-4o-mini as the judge model. (RaR) methods are trained using rubrics generated by either o3-mini or GPT-4o. Across both domains, the RaR-Implicit method consistently outperforms Simple-Likert and matches or exceeds the performance of Reference-Likert.
@@ 4.1 Datasets
We investigate the utility of rubrics as rewards across two reasoning domains medicine and science.
  • RaR‑Medical‑20k: We curate a set of 20k prompts from medicine‑related subsets of medical‑o1‑reasoning‑SFT Chen et al. (2024a), the natural_reasoning dataset Yuan et al. (2025), the SCP‑116K dataset Lu et al. (2025), and the GeneralThought‑430K dataset General Reasoning (2025).
  • RaR‑Science‑20k: We curate roughly 20k science prompts aligned with the GPQA Diamond benchmark from the natural_reasoning dataset Yuan et al. (2025), the SCP‑116K dataset Lu et al. (2025), and the GeneralThought‑430K dataset General Reasoning (2025).
Refer Tables 7 and 10 in Appendix for the full topic distributions.
@@ 4.2 Training Details
We train all models using the GRPO algorithm Shao et al. (2024), with Qwen2.5-7B as the base policy. Training is conducted with a batch size of 96, a learning rate of $5\times 10^{-6}$, and a constant learning rate schedule with 10% linear warmup. Full hyperparameter details are provided in Appendix 11. All models are trained on a single compute node equipped with 8 NVIDIA H100 GPUs.
@@
Reward Computation: We use gpt-4o-mini as the judge model to assign rewards $R_{q}$ to the sampled responses. The reward functions are described in Sections 4.3 and 4.4.
Policy Update: The policy is updated using GRPO based on the computed rewards.
4.3 Baselines
We consider compare various baselines and off-the-shelf post-trained models. All rubric-free baselines are with Qwen2.5-7b as the base policy.
For off-the-shelf baselines we evaluate performance on Qwen2.5-7b. We also include the performance of Qwen2.5-7b-Instruct to compare with instruction tuned variant of the base policy.
We describe the post-trained baselines without rubrics below:
SIMPLE-LIKERT:
LLM-judge is asked to output a Likert Score from 1-10 (normalized to 0-1) for each response-prompt pair.
@@ 4.3 Baselines
The judge model compares the generated response against a high-quality reference answer and outputs a Likert score from 1–10, which is then normalized to the $[0,1]$ range. The reward for each (prompt, response, reference) triplet is defined as:
@@
where $x^{*}$ denotes the reference answer, either written by a human or generated by a stronger LLM.
@@ 4.4 Rubric-guided Methods
PREDEFINED-RaR:
This method uses a fixed set of generic rubrics (not prompt-specific, as described in Appendix C) and computes the reward via an explicit weighted sum, following Equation 1. All rubric criteria are weighted uniformly.
@@
This variant also employs explicit aggregation (1) but uses prompt-specific rubrics generated by the method in Section 3. Since the generated rubrics include categorical labels, we assign numerical weights using a simple default scheme: {"Essential": 1.0, "Important": 0.7, "Optional": 0.3, "Pitfall": 0.8}, with higher penalty applied to violated pitfalls.  [^2] Pitfall criteria are phrased in positive form (e.g., “The response avoids misinformation”), so satisfying them contributes positively to the score. If a pitfall is not satisfied, the corresponding reward is reduced or penalized.
@@
This variant also leverages prompt-specific rubrics; however, instead of aggregating criterion scores explicitly, it prompts the judge model to evaluate the response holistically and assign a single Likert rating (1–10). This rating is then normalized to the $[0,1]$ range:
$R_{\text{implicit}}(q,x)=\mathrm{Norm}(\text{LikertScore}(q,x,R_{q}))$
where $R_{q}$ denotes the list of rubrics associated with prompt $q$.

### v2

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has proven effective for complex reasoning tasks with clear correctness signals such as math and coding. However, extending it to real-world reasoning tasks is challenging, as evaluation depends on nuanced, multi-criteria judgments rather than binary correctness. Instance-specific rubrics have recently been used in evaluation benchmarks to capture such judgments, but their potential as reward signals for on-policy post-training remains underexplored. We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning method that extends RLVR beyond verifiable domains by using rubric-based feedback. Across both medical and science domains, we evaluate multiple strategies for aggregating rubric feedback into rewards. The best RaR variant achieves relative improvements of up to 31% on HealthBench and 7% on GPQA-Diamond over popular LLM-as-judge baselines that rely on direct Likert-based rewards. These results demonstrate that RaR-trained policies adapt well to diverse evaluation formats, performing strongly on both rubric-based and multiple-choice tasks. Moreover, we find that using rubrics as structured reward signals yields better alignment for smaller judges and reduces performance variance across judge scales.
# anisha.gunjal@scale.com
@@ 1 Introduction
Reinforcement Learning with Verifiable Rewards (RLVR) has enabled large language models to elicit complex reasoning on tasks with clear verifiable outcomes. This is especially effective in domains like math and code, where reward models can be replaced by scoring functions or test cases that automatically verify correctness [19, 11, 6]. However, extending RLVR to unstructured, real-world reasoning is challenging because such tasks lack easily verifiable answers. A common workaround is to use preference-based reward models, but they tend to overfit superficial artifacts (e.g. response length, formatting quirks, annotator biases) [29, 34, 4, 36, 10] and require large volumes of pairwise comparisons [25]. Instance-specific rubrics have recently emerged for nuanced evaluation in expert domains [2], yet their application in on-policy training for expert-level reasoning is largely unexplored.
To address this gap, we explore a paradigm shift that introduces a middle ground between the simplicity of verifiable rewards and the expressiveness of preference rankings, which often come with human artifacts and operational overhead. We introduce Rubrics as Rewards (RaR), a framework for on-policy Reinforcement Learning that uses structured criteria or rubrics as the core reward mechanism. Rather than using rubrics only for evaluation [2, 30], we treat them as checklist-style supervision that produces reward signals for on-policy RL. Each rubric is composed of modular, interpretable subgoals that provide automatable feedback aligned with expert intent. By decomposing “what makes a good response” into tangible, human-interpretable criteria, rubrics offer a middle ground between binary correctness signals and coarse preference rankings.
@@
Previous works train generative reward models that learn to evaluate reasoning or final outputs with interpretable scores [5, 35, 1, 12], and some have even used a model’s internal confidence estimates as a proxy for reward [39]. More recent efforts have extended verifiable datasets beyond STEM domains, broadening the applicability of RLVR methods to a wider range of tasks [32, 23]. Yet a general-purpose approach for specifying reliable reward signals remains elusive, particularly in tasks without a single ground truth where both subjective and objective criteria must be considered. In contrast, we treat rubrics as instance-specific, reusable reward functions. Once generated, rubrics provide interpretable and automatable supervision that can be applied consistently across new rollouts, offering a scalable and transparent alternative to opaque reward modeling in on-policy learning.
Recent concurrent works explore checklists and principled rubric criteria for preference tuning and LLM safety [8, 33, 7], highlighting a growing trend toward structured supervision. In contrast, we convert rubrics into reward functions for on-policy RL, targeting expert reasoning and applied real-world domains. This closes the rubric-to-learning loop and improves performance on both rubric-guided evaluations and tasks with verifiable answers. Figure 1 illustrates our framework.
Our key contributions are as follows: (i) We introduce Rubrics as Rewards (RaR), an on-policy reinforcement learning framework that uses checklist-style rubrics for multi-criteria supervision in reasoning and real-world domains. (ii) We synthesize instance-specific rubrics for medicine and science and release the corresponding training sets, RaR-Medicine  [^1] RaR-Medicine: https://huggingface.co/datasets/anisha2102/RaR-Medicine and RaR-Science  [^2] RaR-Science: https://huggingface.co/datasets/anisha2102/RaR-Science. (iii) RaR-trained models consistently outperform strong baselines and yield a stable, generalizable training signal, with gains on both rubric-scored and verifiable multiple-choice evaluation settings. (iv) Our results demonstrate that rubric-based rewards provide stable supervision across judge sizes, helping smaller models align effectively with human preferences and maintaining robust evaluation performance from small to large judges.
@@ 2.1 Problem Formulation
Let $x$ denote an input prompt and $\hat{y}\sim\pi_{\theta}(\cdot\mid x)$ be a sampled response from a model parameterized by $\theta$. In domains without single ground-truth answers or automatic correctness signals, we define a structured reward function using instance-specific rubric criteria.
Each prompt $x$ is associated with a set of $k$ rubric items $\{(w_{j},c_{j})\}_{j=1}^{k}$, where $w_{j}\in\mathbb{R}$ denotes the weight of criterion $j$, and $c_{j}:(x,\hat{y})\mapsto\{0,1\}$ is a binary correctness function that indicates whether the response $\hat{y}$ satisfies that criterion given the prompt.
2.2 Reward Aggregation Strategies
We investigate two complementary approaches for combining rubric feedback into scalar rewards:
Explicit Aggregation.
Each criterion is independently evaluated using an LLM-as-judge, and the final normalized reward is computed as:
@@ 2.2 Reward Aggregation Strategies
Normalization makes rewards comparable across prompts that differ in rubric count or weights. Although we use binary checks for $c_{j}$ in our experiments, the formulation can be extended to continuous-valued scores.
Implicit Aggregation.
All rubric criteria along with categorical weights are passed to an LLM-as-judge, delegating the aggregation to the model itself to produce a single scalar reward:
$r_{\text{implicit}}(x,\hat{y})=f_{\phi}(x,\hat{y},\{d_{j}\}_{j=1}^{k})$ (2)
Here, $f_{\phi}$ denotes an LLM-based judge that takes the prompt $x$, the response $\hat{y}$, and the set of rubric criteria $\{d_{j}\}$ as input. This formulation allows the model to compute a holistic reward score directly, avoiding the need to manually tune rubric weights.
The prompts used for each method are detailed in Appendix A.6.
@@ 2.3 Generalization of RLVR with Rubrics as Rewards
Rubric-based reward functions thus generalize RLVR by enabling multi-dimensional supervision, flexible weighting across criteria, and the incorporation of both objective and subjective aspects of response quality. This formalization highlights that RLVR can be seen as a restricted instance of rubric-guided RL with a single essential criterion. In contrast, rubric-based rewards further enable structured supervision in settings where correctness is multifaceted and may not be strictly verifiable.
@@ 3.1 Desiderata
A rubric specifies criteria for high-quality responses and provides human-interpretable supervision. We identify four desiderata for effective rubric generation:
Grounded in Expert Guidance.
Rubrics should reflect domain expertise by capturing the essential facts, reasoning steps, and conclusions necessary for correctness. Ideally, this grounding comes from human experts or their high-quality proxies.
Comprehensive Coverage.
Rubrics should span multiple dimensions of response quality, including factual accuracy, logical coherence, completeness, style, and safety. Negative criteria (pitfalls) help identify frequent or high-risk errors that undermine overall quality.
Criterion Importance.
Rubrics should reflect that some dimensions of response quality are more critical than others. For example, factual correctness must outweigh secondary aspects such as stylistic clarity. Assigning weights to criteria ensures this prioritization, whether through simple categorical tags, explicit numeric values, or learned weighting schemes.
Self-Contained Evaluation.
Each rubric item should be independently actionable, allowing either human annotators or automated judges to assess it in isolation without requiring external context or domain-specific knowledge.
3.2 Rubrics Creation
We apply these desiderata to datasets for reasoning tasks in medicine and science. Given the scarcity of human-annotated rubric datasets in these domains, we use LLMs to generate instance-specific rubrics from golden reference answers at scale, enabling the study of structured rewards without costly human annotation.
For each prompt, an LLM generates a rubric of 7–20 self-contained items. Each item is assigned both a numeric and a categorical weight reflecting its relative importance. While numeric weights provide fine-grained prioritization, in our experiments we adopt categorical labels (Essential, Important, Optional, Pitfall) for ease of implementation and interpretability in controlled settings. The resulting rubrics are then used directly as reward functions through either explicit aggregation (Eq. 1) or implicit aggregation (Sec. 2.2).
In practice, we generate rubrics using OpenAI’s o3-mini and GPT-4o [24, 15, 14], conditioning generation on reference answers from the underlying datasets to approximate expert grounding. The resulting collections—RaR-Medicine and RaR-Science—are released for public use. These rubric sets supervise smaller policies under GRPO using both explicit and implicit reward aggregation.
@@ 4.1 Datasets
We investigate the utility of rubrics as rewards across two reasoning domains, medicine and science.
  • RaR-Medicine: A dataset of 20k prompts drawn from diverse medical reasoning sources, including medical-o1-reasoning-SFT [3], natural_reasoning [37], SCP-116K [22], and GeneralThought-430K [9]. Instance-specific rubrics for this dataset are generated with GPT-4o (see Appendix A.1).
  • RaR-Science: A dataset of $\sim$20k prompts curated to align with GPQA-Diamond categories. Prompts are sourced from natural_reasoning [37], SCP-116K [22], and GeneralThought-430K [9], covering a broad range of scientific reasoning tasks (Appendix A.2). Rubrics for this dataset are synthesized with o3-mini.
@@ 4.2 Training Details
We conduct all experiments using on-policy reinforcement learning with the GRPO algorithm [28], taking Qwen2.5-7B as the base policy. Models are trained with a batch size of 96, a learning rate of $5\times 10^{-6}$, and a constant schedule with 10% linear warmup. Complete hyperparameter settings are listed in Appendix 11. Training runs are executed on a single compute node equipped with 8 NVIDIA H100 GPUs.
@@
Reward Computation with Rubrics: We use gpt-4o-mini as the judge model to assign rewards $R_{q}$ to the sampled responses. We experiment with various reward computation and aggregations strategies further described in Sections 4.3 and 4.4.
Policy Update: The policy weights are updated using GRPO based on the computed rewards.
4.3 Rubric-Free Baselines
We consider various rubric-free baselines and off-the-shelf post-trained models. Rubric-free baselines are trained with Qwen2.5-7B as the base policy.
OFF-THE-SHELF:
For off-the-shelf baselines we evaluate performance on Qwen2.5-7B. We also include the performance of Qwen2.5-7B-Instruct to compare with instruction-tuned variant of the base policy.
DIRECT-LIKERT:
An LLM-as-judge provides a direct assessment for each response–prompt pair on a 1–10 Likert scale [40, 17], normalized to $[0,1]$. The resulting score is used directly as the reward signal for training.
@@ 4.3 Rubric-Free Baselines
An LLM-as-judge compares the generated response against a reference answer (written by experts or stronger LLMs) and assigns a 1–10 Likert score [40], normalized to $[0,1]$. This reference-guided score is used as the reward signal for policy updates. The reward for each (prompt, response, reference) triplet is defined as:
@@
where $x^{*}$ denotes the reference answer.
@@ 4.4 Rubric-guided Methods
RaR-PREDEFINED:
This method uses a fixed set of generic rubrics for all prompts (e.g. response is concise, response contains correct information). It employs the Explicit Aggregation method (Equation 1) with all criteria weighted uniformly (see Appendix A.5).
@@
This variant also uses Explicit Aggregation using a weighted sum (Equation 1) but applies it to instance-specific rubrics from Section 3. We manually assign numerical weights based on the generated categorical labels: {"Essential": 1.0, "Important": 0.7, "Optional": 0.3, "Pitfall": 0.9} [^3] Pitfall criteria are phrased in positive form (e.g., “The response avoids misinformation”), so satisfying them contributes positively to the score. If a pitfall is not satisfied, the corresponding reward is reduced or penalized..
@@
This variant uses the Implicit Aggregation method (Equation 2). It leverages prompt-specific rubrics, where a judge model evaluates the response as a whole to assign a single Likert rating (1–10), avoiding the need for hand-tuned weights. The reward is normalized to the $[0,1]$ range during training.